Thirty pounds of mixed nuts sells for $6.10 per pound. The mixture is obtained from two kinds of nuts, walnuts priced at $6.50 per pound and peanuts at $5.90 per pound. How many pounds of each variety of nut are used in the mixture?
Answer by jorel1380(3719)

30x6.10=183.00
p+w=30
5.90p+6.50w=183
590p+650w=18300
590p+590w=17700
60w=600
w=10
p=20
You need 20 lbs. of peanuts and 10 lbs. of cashews for the mix..